Compute for the pH and pOH of a solution having a 0.3 M H3PO4 and 0.1 M NaHPO4 buffer solution


1
Expert's answer
2021-10-05T01:30:33-0400

pH=0.3+0.1/22.4l/mol*100=1.7

pOH=14-1.7=12.3